### v1

# CEval: A Benchmark for Evaluating Counterfactual Text Generation

Abstract
Counterfactual text generation aims to minimally change a text, such that it is classified differently. Judging advancements in method development for counterfactual text generation is hindered by a non-uniform usage of data sets and metrics in related work. We propose CEval, a benchmark for comparing counterfactual text generation methods. CEval unifies counterfactual and text quality metrics, includes common counterfactual datasets with human annotations, standard baselines (MICE, GDBA, CREST) and the open-source language model LLAMA-2. Our experiments found no perfect method for generating counterfactual text. Methods that excel at counterfactual metrics often produce lower-quality text while LLMs with simple prompts generate high-quality text but struggle with counterfactual criteria. By making CEval available as an open-source Python library, we encourage the community to contribute more methods and maintain consistent evaluation in future work. [^1] The source code is included with the paper submission and will be publicly accessible upon acceptance.
CEval: A Benchmark for Evaluating Counterfactual Text Generation
Van Bach Nguyen University of Marburg / Germany vanbach.nguyen@uni-marburg.de                        Jörg Schlötterer University of Marburg / Germany University of Mannheim / Germany joerg.schloetterer@uni-marburg.de
Christin Seifert University of Marburg / Germany christin.seifer@uni-marburg.de
1 Introduction
The growing popularity of AI and increasingly complex “black-box” models triggered a critical need for interpretability. As Miller (2019) highlights, explanations are often counterfactual, seeking to understand why event P occurred instead of alternative Q. Ideally, explanations should demonstrate how minimal changes in an instance could have led to different outcomes. For example, given a review: The film has funny moments and talented actors, but it feels long. To answer the question why this review has a negative sentiment instead of a positive sentiment, the answer might involve showing a positive example as a counterfactual like: The film has funny moments and talented actors, yet feels a bit long. This example enables us to identify specific words that require change and the necessary modifications to achieve the target sentiment. This motivates counterfactual generation, the task of modifying an instance to produce a different model prediction with minimal change.
Besides explanation Robeer et al. (2021), the NLP community also utilizes counterfactual generation for various purposes such as debugging models Ross et al. (2021), data augmentation Dixit et al. (2022); Chen et al. (2023); Bhattacharjee et al. (2024) or enhancing model robustness Treviso et al. (2023); Wu et al. (2021). However, generating counterfactuals is not straightforward due to the complexity of textual changes, involving replacements, deletions, and insertions. Determining where and how to modify text to alter predictions remains an open issue. Existing research efforts unified evaluation standards and often prioritize quantity over quality in generated counterfactuals. Table 1 illustrates the disparity in datasets, metrics, and baselines across different studies. Consequently, it becomes challenging to choose an optimal method for a specific purpose. This highlights the need for a well-defined benchmark to comprehensively evaluate counterfactual generation methods for textual data. Such a benchmark should establish standard datasets, metrics, and baselines, enabling fair and meaningful comparisons.
Method Dataset Metrics Baseline
MICE Ross et al. (2021) IMDB, Race, Newgroups Flip rate, Fluency, Minimality MICE’s variants
GBDA Guo et al. (2021) AG News, Yelp IMDB, MNLI Accuracy, Cosine Similarity, #Queries TextFooler Jin et al. (2020) BAE Garg and Ramakrishnan (2020) BERT-Attack Li et al. (2020)
CF-GAN Robeer et al. (2021) HATESPEECH, SST-2, SNLI Fidelity Perceptibility Naturalness SEDC Martens and Provost (2014) PWWS+ Ren et al. (2019) Polyjuice Wu et al. (2021) TextFooler Jin et al. (2020)
Polyjuice Wu et al. (2021) IMDB, NLI, QQP Diversity, Closeness GPT-2 Radford et al. (2019) T5 Raffel et al. (2020) RoBERTa Liu et al. (2019)
Table 1: The variations in datasets, metrics, and baselines used across different methods.
This work introduces CEval - a comprehensive benchmark to unify the evaluation of methods that modify text to change a classifier’s prediction. Such methods include contrastive explanations, counterfactual generation, and adversarial attacks. Our benchmark offers multifaceted evaluation: assessing both the “counterfactual-ness” (e.g., label flipping ability) and textual quality (e.g., fluency, grammar, coherence) of counterfactual texts. Additionally, it comprises carefully curated datasets with human annotations, as well as a straightforward baseline generated using a large language model. Moreover, we systematically review state-of-the-art methods in this field and directly compare their performance in our benchmark. We provide an open-source library for both the benchmark and the methods, promoting reproducibility and facilitating further research in this domain. Furthermore, we present a systematic comparison of different methods, highlighting their strengths and weaknesses in generating counterfactual text. We analyze how automatically generated counterfactuals compare to human examples, revealing gaps and opportunities for improvement. We find that counterfactual generation methods often generate text that lacks in quality compared to simple prompt-based LLMs. While the latter may struggle to satisfy counterfactual metrics, they typically exhibit higher text quality. Based on these insights, we suggest exploring combinations of methods as promising directions for further research.
2 Related Work
While terms like “counterfactual” and “contrastive” generation are often used interchangeably in literature Stepin et al. (2021), our work adopts a specific definition. We define counterfactual generation as generating a new instance with different model predictions from the original with minimal changes, encompassing both true counterfactuals and contrastives. This broader category includes counterfactual, contrastive generation, and adversarial attacks. In the past, adversarial attacks focused on changing the label without considering text quality. Recent work like GBDA Guo et al. (2021) focuses on producing adverserial text that is more natural by adding fluency and semantic similarity losses. Hence, we include GBDA in our benchmark. Technically, counterfactual generation methods for text fall into three categories:
Masking and Filling Methods (MF): These methods perform 2 steps: (1) identify important words for masking and use various techniques such as selecting words with the highest gradient or training a separate rationalizer for the masking process. (2) The masked words are then replaced using a pretrained language model with fill-in-blank ability. In step (1), MICE Ross et al. (2021) and AutoCAD Wen et al. (2022) use the gradient of the classifier. DoCoGen Calderon et al. (2022) masks all relevant terms in a domain, while CREST Treviso et al. (2023) trains a separate rationalizer, i.e., SPECTRA Guerreiro and Martins (2021). Then, all of them fine-tune T5 to fill in blanks but train in different ways. Polyjuice Wu et al. (2021) accepts text with masking and fine-tunes a Roberta-based model to fill in blanks using control codes.
Conditional Distribution Methods (CD): Methods like GBDA Guo et al. (2021) and CF-GAN Robeer et al. (2021) train a conditional distribution for counterfactuals. The counterfactuals are obtained by sampling from this distribution based on a target label.
Counterfactual Generation with Large Language Models: Recently, there has been a trend towards using Language Models (LLMs) for counterfactual generation. Approaches like CORE Dixit et al. (2022), DISCO Chen et al. (2023) and FLARE Bhattacharjee et al. (2024) optimize prompts fed into LLMs to generate the desired counterfactuals. This trend is driven by the versatile capabilities of LLMs in various tasks Maynez et al. (2023).
Despite the diverse approaches employed in generating counterfactuals across various studies, the common objective remains to generate high-quality counterfactuals. However, these studies employ different metrics, baselines, and datasets, as illustrated in Table 1. Therefore, given the rapid growth of approaches in this field, establishing a unified evaluation standard becomes paramount. Existing benchmarks for counterfactual generation Pawelczyk et al. (2021); Moreira et al. (2022) focus exclusively on tabular data with properties that are orthogonal to text (e.g., continuous value ranges). Hence, we introduce CEval to fill this gap and provide a standard evaluation framework specifically tailored to textual counterfactual generation. Our benchmark unifies metrics of both, counterfactual criteria and text quality assessment, including datasets with human annotations and a simple baseline from a large language model.
3 Benchmark Design
We focus on counterfactual generation for textual data, which involves editing existing text with minimal modifications to produce new text that increases the probability of a predefined target label with respect to a black-box model. This process aims to generate a counterfactual, denoted as $x^{\prime}$, that alters the model’s predictions compared to the original text $x$.
Formally, given a fixed classifier $f$ and a dataset with $N$ samples $(x_{1},x_{2},\ldots,x_{N})$, $x=(z_{1},z_{2},\ldots,z_{n})$ represents a sequence of $n$ tokens. The original prediction is denoted as $f(x)=y$, while the counterfactual prediction is $y^{\prime}\neq y$. The counterfactual generation process is represented by a method $e:(z_{1},\ldots,z_{n})\mapsto(z^{\prime}_{1},\ldots,z^{\prime}_{m})$, ensuring that $f(e(x))=y^{\prime}$. The resulting counterfactual example is $x^{\prime}=(z^{\prime}_{1},\ldots,z^{\prime}_{m})$ with $m$ tokens.
A valid counterfactual instance should satisfy the following criteria Molnar (2022):
Predictive Probability: A counterfactual instance $x^{\prime}$ should closely produce the predefined prediction $y^{\prime}$. In other words, the counterfactual text should effectively lead to the desired target label.
Textual Similarity: A counterfactual $x^{\prime}$ should maintain as much similarity as possible to the original instance $x$ in terms of text distance. This ensures that the generated text remains coherent and contextually aligned with the original.
Likelihood in Feature Space: A valid counterfactual instance should exhibit feature values that resemble real-world text, indicating that $x^{\prime}$ remains close to a common distribution for text. This distribution can be represented by pretrained large language models trained on massive text corpora. This criterion ensures that the generated text is plausible, realistic and consistent with typical language patterns.
Diversity: There are multiple ways to modify the input to reach the desired label. A good counterfactual method should provide various options for changing a text instance to obtain the target label.
3.1 Metrics
In CEval, we use two types of metrics: counterfactual metrics, which reflect the counterfactual criteria outlined above, and textual quality metrics, which assess the quality of the generated text, irrespective of its counterfactual properties.
3.1.1 Counterfactual metrics
Flip Rate (FR): measures how effectively a method can change labels of instances with respect to a pretrained classifier. This metric represents the binary case of the Predictive Probability criterion, determining whether the label changed or not and is commonly used in the literature Treviso et al. (2023); Ross et al. (2021). FR is defined as the percentage of generated instances where the labels are flipped over the total number of instances $N$ Bhattacharjee et al. (2024):
$FR=\frac{1}{N}\sum_{i=1}^{N}\mathbbm{1}[f(x_{i})\neq f(x^{\prime}_{i})]$
Probability Change ($\Delta\textbf{P}$): While the flip rate offers a binary assesment of Predictive Probability, it does not fully capture how closely a counterfactual instance aligns with the desired prediction. Some instances may get really close to the target prediction but still fail to flip the label. To address this limitation, we propose the Probability Change ($\Delta$P) metric, which quantifies the difference between the probability of the target label for the original instance $x$ and the probability of the target label for the contrasting instance $x^{\prime}$.
$\Delta P=\frac{1}{N}\sum_{i=1}^{N}(P(y^{\prime}_{i}|x_{i},f)-P(y^{\prime}_{i}|x^{\prime}_{i},f))$
Here, $P(\cdot)$ represents the probability of having the prediction $y$ given the instance $x$ with respect to the classifier $f$. The expression denotes the difference in the probabilities of the prediction $y^{\prime}$ for instances $x$ and $x^{\prime}$ according to the classifier $f$. Token Distance (TD): To measure Textual Similarity, a common metric employed in literature Ross et al. (2021); Treviso et al. (2023) is the token-level distance $d(x,x^{\prime})$, where $d$ represents any edit distance metric. In this work, we use the Levenshtein distance for $d$ to quantify the token-distance between the original instance $x$ and the counterfactual $x^{\prime}$. This choice is motivated by the Levenshtein distance’s ability to capture all type of edits (insertions, deletions, or substitutions) and also its widespread use in related work Ross et al. (2021); Treviso et al. (2023):
$TD=\frac{1}{N}\sum_{i=1}^{N}\frac{{d(x_{i},x^{\prime}_{i}})}{|x_{i}|}$
Perplexity (PPL): To ensure that the generated text is plausible, realistic, and follows a natural text distribution, we leverage perplexity from GPT-2 because of its effectiveness in capturing such distributions. Radford et al. (2019) [^2] While GPT-2 is used for simplicity in this study, any other LLM can be substituted as long as it demonstrates strong text generation capabilities
$PPL(x)=\exp\left\{-\frac{1}{n}\sum_{i=1}^{n}\log p_{\theta}(z_{i}\mid z_{<i})\right\}$
Diversity (Div): We quantify diversity by measuring the token distance between pairwise generated instances. Given two counterfactuals, $x^{\prime 1}$ and $x^{\prime 2}$, for the same instance $x$, diversity is defined as the mean pairwise distance between the sets of counterfactuals:
$Div=\frac{1}{N}\sum_{i=1}^{N}\frac{d(x^{\prime 1}_{i},x^{\prime 2}_{i})}{|x_{i}|}$
Here, $d(x^{\prime 1}_{i},x^{\prime 2}_{i})$ represents the Levenshtein distance between the corresponding tokens of the two counterfactuals for the $i$-th instance, and $|x_{i}|$ is the length of the original instance $x_{i}$.
3.1.2 Text Quality Metrics
In addition to counterfactual evaluation metrics, it is essential to consider the quality of the generated text. Text quality metrics are designed to evaluate specific aspects of textual content. Following Chiang and Lee (2023), key text quality metrics include:
Fluency: Relates to the natural and fluent flow of the generated text. A fluently constructed counterfactual instance should exhibit coherence in its structure and readability, ensuring that the transformed text reads naturally and seamlessly.
Cohesiveness: Focuses on the coherence and logical flow of ideas within the generated text. A cohesive counterfactual instance should deliver a cohesive narrative, with ideas and sentences seamlessly transitioning from one to another, contributing to an overall consistent and well-structured piece.
Likability: This metric focuses on how pleasant and likable the generated text is perceived to be by readers. It considers factors like tone, style, and the overall user experience. A counterfactual instance that scores high on likability is expected to be more engaging and appealing to readers.
Grammar: Evaluates the syntactical correctness and grammatical accuracy of the generated text. A counterfactual instance should follow grammatical rules, ensuring that sentences are properly constructed, punctuation is used correctly, and overall grammar is of a high standard.
The given textual metrics play a crucial role in evaluating generated text comprehensively, offering insights into various quality aspects. When combined with traditional counterfactual metrics, they provide a detailed understanding of both effectiveness and linguistic quality. Chiang and Lee (2023) demonstrated that language models (LLMs) can serve as an alternative method to assess text quality using these metrics. In this work, we use Mistral-7B-instruct Jiang et al. (2023) to evaluate the above textual quality metrics on a scale ranging from 1 to 5, as outlined in prior work Chiang and Lee (2023).
3.2 Datasets
To evaluate against human-generated counterfactuals, we selected datasets with contrast sets generated by crowd-sourcing Kaushik et al. (2020) or human experts Gardner et al. (2020). We chose two common benchmark datasets capturing two different NLP tasks: sentiment analysis IMDB Maas et al. (2011) and natural language inference (NLI) on SNLI Bowman et al. (2015). Further datasets with pre-trained classifiers can be easily added to the benchmark.
IMDB contains movie reviews in a diverse range of language styles, sentiments, and topics ensuring that models trained on this dataset are exposed to various linguistic patterns and expressions. Following (Maynez et al., 2023), we select the 488 instances annotated by humans. The dataset is evenly balanced (243 negative and 245 positive instances). We use a pre-trained BERT model [^3] https://huggingface.co/textattack/bert-base-uncased-imdb from TextAttack Morris et al. (2020) as target classifier with an accuracy of 89%. The counterfactual text generation task is to edit the review instances, whether positive or negative, to change the prediction of the classifier on that review with minimal edits. IMDB contains both, contrast examples generated by human experts Gardner et al. (2020) and by crowdworkers Kaushik et al. (2020).
SNLI Bowman et al. (2015) consists of pairs of sentences with a label indicating whether they entail, contradict, or are neutral to each other. NLI requires models to understand the semantic relationships between sentences, which is important for various downstream NLP applications Camburu et al. (2018). Again, we a pre-trained BERT model [^4] https://huggingface.co/textattack/bert-base-uncased-snli from TextAttack Morris et al. (2020) as target classifier for the NLI task with an accuracy of 90%. The counterfactual text generation task is to modify either the premise or hypothesis in the sentence pairs to flip the classifier’s label.
4 Method Selection
In this section, we review state-of-the-art methods for generating counterfactuals, carefully selected from a diverse set of methods found in the literature. To narrow down our comparison in the benchmark, we establish specific properties for method selection. Our criteria aim to make the evaluation comprehensive with minimal effort. We consider not only methods explicitly designed for counterfactual generation but also those with a shared goal of producing high-quality text with respect to a target label. Consequently, we incorporate approaches from the domain of adversarial attacks, expanding the scope beyond traditional counterfactual and contrastive generation methods.
To select methods for the benchmark, we establish the following criteria:
Type of Methods: We categorize methods into Masking and Filling, Conditional Distribution, or Counterfactual Generation with Language Models as mentioned in Section 2. Each category should be covered by a representative method.
Classifier Accessibility: We consider whether the method has access the classifier during the counterfactual generation process.
Reproducibility: We check whether the method provides reproducible source code, ensuring transparency and replicability.
Problem Agnostic: We evaluate whether the method is versatile enough to encompass various types of classification problems.
After filtering based on these criteria (detailed in the Appendix, Table 5), we determine the final set of methods as follows:
MICE Ross et al. (2021) is a contrastive explanation generation method. The idea is, given a predictor and a dataset, MICE trains an editor to generate tokens to fill masked tokens, such as the final text flip the original label. To determine which tokens and how many tokens to mask, the method uses the list of tokens with the highest gradients contributing to the predictions. It employs binary search to find the minimum number of tokens that need to be masked. This method requires access to the classifier to verify the label internally, representing a counterfactual generation method.
IMDB SNLI
LLAMA-2 MICE GBDA CREST Crowd Expert LLAMA-2 MICE GBDA CREST Crowd
CF Metrics Flip Rate $\uparrow$ 0.66 0.99 0.97 0.71 0.81 0.84 0.34 0.85 0.94 0.39 0.75
$\Delta$Probability $\uparrow$ 0.65 0.91 0.96 0.70 0.80 0.83 0.27 0.65 0.86 0.10 0.64
Perplexity $\downarrow$ 41.3 62.1 84.7 44.0 55.9 47.2 54.3 160 143 60.9 72.1
Distance $\downarrow$ 0.23 0.20 0.25 0.33 0.16 0.14 0.58 0.47 0.47 0.32 0.41
Diversity $\uparrow$ 0.24 0.25 0.26 0.43 0.24 0.25 - - - - -
Text Quality Grammar $\uparrow$ 4.80 4.35 3.47 3.89 4.71 4.79 4.83 4.63 3.80 4.21 4.79
Cohesiveness $\uparrow$ 4.63 4.30 3.61 3.53 4.71 4.62 3.67 3.30 2.08 2.83 3.63
Likability $\uparrow$ 2.82 3.12 3.07 2.65 3.12 2.96 3.81 3.31 2.25 3.19 3.52
Fluency $\uparrow$ 3.83 3.58 3.36 3.20 3.81 3.73 4.71 4.60 3.90 4.20 4.67
Average $\uparrow$ 4.02 3.84 3.38 3.32 4.09 4.02 4.25 3.96 3.01 3.61 4.15
Table 2: Results with counterfactual (CF) and text qualily metrics on IMDB and SNLI [^6] We omit diversity, as SNLI has no human expert labels.. Average denotes mean of text quality metrics, each scored on a scale 1-5 following Chiang and Lee (2023)).
GBDA Guo et al. (2021) is a gradient-based adversarial attack, a white-box attack that leverages a novel adversarial distribution formulation, enabling end-to-end optimization of the adversarial loss and fluency constraints with gradient descent. This method requires the use of a model as a classifier, which can be the same as the pretrained model or different, depending on whether access to the original classifier is needed. This method represents the adversarial attack domain.
CREST Treviso et al. (2023) has a similar approach to MICE as they train a rationalizer using SPECTRA Guerreiro and Martins (2021) to find the optimal masks instead of using the highest gradient tokens. Then, they apply filling with T5 following MICE. Given the popularity of Mask and Filling type, we choose another method for a more comprehensive comparison.
LLAMA-2 Touvron et al. (2023): Large Language Models have shown impressive performance on many tasks with only simple prompts Srivastava et al. (2023). Therefore, in this study, we use LLAMA-2 with simple one-shot learning as a baseline for comparing other methods. This choice is made in contrast to other works that used closed-source LLMs.
5 Results
We evaluate the performance of all compared methods (LLAMA-2, MICE, GBDA, CREST) against human crowd-sourced and human expert generations. It is important to note that MICE and GBDA have access to the prediction model during generation, while CREST leverages a pre-trained T5 model and transfers its findings to the target BERT model. In contrast, LLAMA-2 and both human evaluation groups (crowd and expert) generate counterfactual examples solely based on the provided text and prompt.
Method Text Predicted label
Original If you haven’t seen this, it’s terrible. It is pure trash. I saw this about 17 years ago, and I’m still screwed up from it. Negative (99.94%)
LLAMA-2 If you haven’t seen this, it’s terrible a masterpiece. It is pure trash brilliance. I saw this about 17 years ago, and I’m still in screwed up awe from it. Positive (99.95%)
MICE If you haven’t seen this, it’s terrible pretty. It is pure trash genius. I saw this about 17 years ago, and I’m still screwed up from it. Positive (99.92%)
GBDA If you haven’t seen this, it’s terrible complicated. It is pure trash the magic. I saw it about 17 30 years ago, and I’m still screwed reeling up from it. Positive (99.65%)
CREST If you haven’t seen this movie, it’s terrible definitely worth seeing. It is pure trash ’s great. I saw it about 17 years ago, and I’m still screwed up from it. Positive (99.89%)
Expert If you haven’t seen this, it’s terrible incredible. It is pure trash gold. I saw this about 17 years ago, and I’m still screwed pumped up from it. Positive (99.89%)
Crowd If you haven’t seen this, it’s terrible incredible. It is pure trash gold. I saw this about 17 years ago, and I’m still screwed up hype about it. Positive (99.52%)
Table 3: Example where all methods successfully flipped the label
We start with an illustrative example: Table 3 shows the shortest example in the IMDB dataset where all methods, including human edits, manage to change the predicted label of the original sentence. This example provides a sneak preview on the methods’ varying properties and issues that the quantitative results reveal. For this simple instance, all methods and human groups agree on replacing negative words like terrible and trash with positive words, even though they differ in their choice of positive words. GDBA is the only exception, its replacements do not always convey a positive sentiment, which reduces text quality. Similarly, MICE and CREST fail to detect the negative phrase screwed up, which renders the text less cohesive and fluent then the text generated by LLAMA-2 and humans, who adapt this negative phrase as well. Besides correctly identifying important words, GDBA also replaces irrelevant words like 17 30, resulting in a larger edit distance. For a more complex example with higher variation of edits and generated text, see Table 6 in the Appendix.
There is no single best method.
The results in Table 6 show that no method outperforms all the others, even when we consider only a single dataset. Expectably, methods with access to the target classifier (MICE & GDBA) are most successful in flipping the label, surpassing even both human groups. However, this success comes at the cost of lower text quality and high perplexity: MICE & GDBA generate “unnatural” text. The fact that humans are less successful in flipping the label indicates limitations within the target classifier, as we would expect a perfect (100%) flip rate for human generation, due to its “gold standard” status. Such potential issues are consistent with prior studies Kaushik et al. (2020); Gardner et al. (2020).
Models and Humans are similarly diverse.
Except for CREST, all methods obtain similar (low) diversity values that are also similar to human diversity values. We attribute these low values to the fact that minimal changes limit the amount of variation. Consequently, CREST has a larger distance to the original text. These observations indicate that if a diverse set of generated text is required, this set should be constructed from the output of different methods.
Counterfactual methods suffer reduced text quality.
In terms of text quality metrics, text generation by humans and LLAMA-2 consistently outperforms other automated methods on most metrics across both datasets. This discrepancy indicates that a focus on counterfactual metrics comes at the cost of reduced text quality. While in particular methods with access to the target classifier (MICE & GDBA) excel at flip rate, they lack in grammar and cohesiveness. Interestingly, LLMs as evaluation proxy seem to prefer the output of another LLM over human generated text. In particular on the SNLI dataset, our evaluation proxy Mistral-7B-instruct consistently prefers LLAMA-2 over human generated text across all text quality metrics. Similarly, GPT-2 as model for measuring perplexity favors LLAMA-2, both on IMDB and SNLI.
Probability changes are mostly bimodal.
Interestingly, MICE has the highest flip rate (FR), but not the largest change in target label probability change ($\Delta$P) on the IMDB dataset. We observe a similar pattern when comparing LLAMA-2 and CREST on the SNLI dataset. CREST has a higher FR, despite LLAMA-2 inducing a larger $\Delta$P.
Figure 1: Histogram of target label probabilities of all methods on the IMDB dataset, including original text and human groups.
The histogram in Figure 1 shows that only MICE generates a significant amount of instances that are close to the decision boundary. All others, including human groups and the original text, exhibit a bimodal pattern with narrow peaks at the two extremes. While the imperfect FR of human groups indicated limitations in the target classifier, the histogram pattern may indicate the source of those limitations: This pattern points to a poorly calibrated, overconfident target classifier, a common issue in today’s deep learning architectures Guo et al. (2017).
LLAMA-2 MICE GBDA CREST Crowd
E N C E N C E N C E N C E N C
Grammar 4.89 4.94 4.57 4.79 4.67 4.41 4.12 4.00 3.50 4.40 3.84 3.35 4.84 4.84 4.70
Cohesiveness 4.29 4.12 2.01 4.26 3.47 2.31 2.86 2.33 1.58 3.19 1.97 1.55 4.08 3.94 3.06
Likability 4.29 4.06 2.65 4.18 3.41 2.58 2.90 2.39 1.83 3.56 2.31 1.87 3.79 3.72 3.26
Fluency 4.99 4.86 4.38 4.90 4.67 4.38 4.61 4.07 3.56 4.43 3.73 3.13 4.95 4.83 4.30
Average 4.61 4.50 3.40 4.53 4.06 3.42 3.62 3.20 2.62 3.90 2.96 2.48 4.42 4.33 3.83
Table 4: Textual quality metrics to verify the LLMs evaluation. E: Entailment, N: Neutral, C: Contradiction
We plan to investigate the impact of a calibrated classifier in future work.
Generated texts exhibit substantial differences.
MICE has the lowest distance among automated methods [^7] In Table 6 we report distance only for true counterfactuals. on the IMDB dataset, but still falls short of human performance. Except for CREST, which has the highest distance on IMDB and lowest on SNLI, distances of all other methods are in a similar range within the same dataset. Despite similar distance ranges, edits might be qualitatively different. To investigate the similarity of edits by different methods, we calculated the average pairwise distance between all generated examples on the IMDB dataset, regardless of label flip success. The results are visualized in Figure 2.
Figure 2: Avg. pairwise Levenshtein distance on IMDB.
Human crowd-sourced and expert edits are highly similar, indicating substantial overlap in their modifications. LLAMA-2 generated text is closest to human edits, which makes it the most promising candidate to serve as proxy for human-generated counterfactuals. Since LLAMA-2’s edits are so close to human edits, we also attribute part of the flip rate failure to the target classifier, rather than to LLAMA-2 (cf. discussion of expected perfect human flip rate above). GBDA and CREST have the largest distance to all other methods (including the original text) and to each other, i.e., their edits are largely distinct. This substantial difference in generated texts suggests that robustness analyses of the target classifier should always be conducted with multiple methods.
LLMs are a valid evaluation proxy for text quality evaluation.
Previous studies Chiang and Lee (2023); Liu et al. (2023) showed that LLMs can be an alternative to human evaluation. In our results, we observe a tendency of LLMs to prefer the output of other LLMs, both in terms of text quality and perplexity, albeit differences are subtle. To validate Mistral-7B-instruct as text quality evaluation proxy, we compare textual quality metrics of SNLI across two labels: contradiction and entailment. We hypothesize that entailment pairs have higher cohesiveness, fluency, and likability compared to contradicting pairs, as entailment implies a relationship between the two sentences. The results in Table 4 show that entailment pairs score significantly higher in text quality than contradicting pairs, especially in cohesiveness and fluency. We observe consistently higher scores across all methods and human generated text. These findings confirm our hypothesis and we see the same pattern in neutral pairs. As expected, differences in neutral pairs are less pronounced with average values from 2.96 (CREST) to 4.50 (LLAMA-2). [^8] All experiments were conducted with a single NVIDIA A6000 GPU. A full benchmark iteration from counterfactual text generation to evaluation including all methods took approximately 10 days.
6 Conclusion
We propose CEval to raise awareness and suggest a standard way to evaluate counterfactual generation in text data, an area gaining traction in the NLP community. Given the nature of textual data, we emphasize the importance of both, counterfactual metrics and text quality. Our benchmark not only fosters standardized comparisons, but also analyzing strengths and weaknesses of individual methods. Our instantiation of the benchmark with representative methods revealed that counterfactual methods excel in counterfactual metrics, but suffer degraded text quality. In turn, LLMs produce high quality text, but struggle to reliably flip the label. Combining the strengths of both approaches opens avenues for future research, e.g., by guiding an LLM’s output with supervision from the target classifier. Low diversity within methods and high diversity across methods highlights the need to conduct robustness analyses of the target classifier with multiple methods. An imperfect flip rate of human generated text (and partly LLMs) can reveal predictive performance issues of the target classifier. The target classifier in our experiments seems to be poorly calibrated and we plan to investigate the impact of well calibrated classifiers in future work. By providing CEval as an open-source Python library, we encourage the community to contribute additional methods and ensuring future work follows the same standards.
Limitations
We employ default hyperparameters for each method and straightforward prompts with LLMs, which may not be optimal for the task at hand and could be furtrher improved by hyperparameter optimization and prompt engineering.
This benchmark solely evaluates the quality of counterfactual text. Further research is required to evaluate the performance of this text in downstream tasks such as classification, data augmentation, or robustness improvement. Additionally, we evaluate the metrics with a single classifier. While this classifier achieves state-of-the-art classification accuracy, our results indicate that it might not be well calibrated. Estimating to which extent our findings can be generalized requires a combination of multiple diverse classifiers in the benchmark and the application in downstream tasks.
Appendix A Method Selection Criteria
Method Type Classifier Access Reproducible code Problem Agnostic
MICE MF
CF-GAN CD
Polyjuice MF
GBDA CD
DISCO LLM
AutoCAD MF
CORE MF
DoCoGen MF
Tailor Ross et al. (2022) MF
CREST MF
GYCMadaan et al. (2021) CD
FLARE LLM
Table 5: Comparison of Methods. Those methods with different types meeting all criteria (highlighted in bold) have been selected for inclusion in the benchmark.
Appendix B Generated Text Comparison Example
Table 6 presents examples where the majority of methods were unsuccessful in altering the original label. While LLAMA-2 and human evaluators both identify nonsensical words within the text, other methods overlook this aspect. In this intricate example, human crowdsource agreement with the human expert is not notably high, as their concurrence is limited to the term nonsensical. However, the human expert’s observations exhibit more alignment with other methods, such as modifying denigrate akin to LLAMA-2, and replacing Sorry or nonsense as observed in MICE.
Method Text Predicted label
Original This movie frequently extrapolates quantum mechanics to justify nonsensical ideas, capped by such statements like "we all create our own reality". Sorry, folks, reality is what true for all of us, not just the credulous. The idea that "anything’s possible" doesn’t hold water on closer examination: if anything’s possible, contrary things are thus possible and so nothing’s possible. This leads to postmodernistic nonsense, which is nothing less than an attempt to denigrate established truths so that all ideas, well-founded and stupid, are equal. To quote sci-fi writer Philip K. Dick, who put it so well, "Reality is that which, when you stop believing in it, doesn’t go away. Negative
LLAMA-2 This movie frequently extrapolates quantum mechanics to justify nonsensicalinspiring ideas, capped by such statements like "we all create our own reality". Sorry, folks, reality is what true for all of us, not just the credulous. The idea that "anything’s possible" doesn’t hold water on closer examination: if anything’s possible, contrary things are thus possible and so nothing’s possible. This leads to postmodernistic nonsense, which is nothing less than an attempt to denigratecelebrate established truths so that all ideas, well-founded and stupid, are equal. To quote sci-fi writer Philip K. Dick, who put it so well, "Reality is that which, when you stop believing in it, doesn’t go away. Negative
MICE This movie frequently extrapolates excellent film has nothing more to say than to condemn quantum mechanics to justifybetray nonsensical ideas, cappedaccompanied by such statements like "we all create our own reality". Sorry,Hey, folks, reality is what true for all of us, not just the credulous. The idea that "anything’s possible" doesn’t hold water on closer examination: if anything’s possible, contrary things are thus possible and so nothing’s possible. This leadsmovie is intended to postmodernistic nonsense, which teach believers that embracing reality is nothing less than an attempt excuse to denigrate established truths so that all ideas, well-founded and stupid, doubtful, are equal. To quote sci-fi writer Philip K. Dick, who put it so well, "Reality is that which, when you stop believing in it, doesn’t go away. Positive
GBDA this movie frequently still extrapolates quantum mechanics experimental depression to justify such nonsensical ideas, capped accompanied by such false statements like like " we all create our own reality " . sorry, folks, reality ". nonetheless, nonetheless, irony iswhattruewhat, for all of us, not just the credulous. the idea that " anything’s possible " doesn’t hold water on closer examination: go away for subjective assumptions : if anything’s possible, contrary everyday things are thus ever possible and so nothing’s everything’s possible. this leads applies to postmodernisticpostmodernist nonsense, authenticity, which is nothing less than an attempt to denigrate established truths cultural reality so that all those ideas, well-founded well - beautiful and stupid, beautiful, are equal. wonderful. to quote sci-fi writer sci - fi critic philip k. dick, who put points it so well, "reality " comedy is that which, when you stop believing in it, yourself, doesn’t go away. Positive
CREST This movie frequently extrapolates quantum mechanics to justify nonsensical ideas, cappedA quantum-sensical thriller, accompanied by such statements films like "we all create our own reality" world". Sorry, folks, thisreality is not what true for all of us, not just the the credulouscredulity.The idea that "anything’s possible" doesn’t hold water on closer-to-end: closer examination: if anything’s possible, contrary things are thus possible and so nothing’s that’s possible. This leads However, there is no less reason to definately postmodernistic nonsense, which is nothing less than an attempt to denigrate established truths characters so that all ideas, the characters, well-founded and stupid, well-meaning, are equal. not. To quote sci-fi writer Philip K. Dick, who put it so well, this film together, "Reality, "Really, is that which, when you stop believing in it, it doesn’t go away. Negative
Expert This movie frequently extrapolates quantum mechanics to justify nonsensical futurist ideas, capped by such inspiring statements like "we all create our own reality". Sorry, Yes, folks, reality is this, what true for all of us, is what we just see, not just the credulous. The idea that "anything’s possible" doesn’t hold water even on closer examination: if anything’s possible, contrary things are thus possible and so nothing’s possible. possible but we’re talking alternate universe. This leads to postmodernistic nonsense, theories, which is are nothing less than an attempt to denigrate elevate established truths so that all ideas, well-founded and stupid, are equal. To quote sci-fi writer Philip K. Dick, who put it so well, "Reality is that which, when you stop believing in it, doesn’t go away. Negative
Crowd This movie frequently extrapolates quantum mechanics to justify nonsensical wise ideas, capped by such statements like "we all create our own reality". Sorry, folks, reality is what true for all of us, not just the credulous. The idea that "anything’s possible" doesn’t hold water on closer examination: if anything’s possible, contrary things are thus possible and so nothing’s possible. This leads to postmodernistic nonsense, which is nothing less than an attempt to denigrate established truths so that all ideas, well-founded and stupid, are equal. To quote sci-fi writer Philip K. Dick, who put it so well, "Reality is that which, when you stop believing in it, doesn’t go away." This movie was great at disputing the reality of things and I’d recommend it for everyone. Negative
Table 6: Example that most methods failed to flip the label